# Group Therapy for Halos: Advancing Halo Mass Estimation for Galaxy Groups

### v1

@@ Appendix B SHMR Fitting Parameters
The calibration of the stellar-to-halo mass relation (SHMR) was performed using a Bayesian Markov Chain Monte Carlo (MCMC) approach, enabling a rigorous exploration of the posterior probability distributions for all model parameters. The adopted functional form for the SHMR is a double power-law,
@@
The resulting posterior distributions and covariances for the four free parameters are shown in Figure B9. The median values and 68% credible intervals for each parameter are:
$A$ $=46.944^{+0.526}_{-0.514},$
$\log M_{A}$ $=10.483^{+0.032}_{-0.030}M_{\odot},$
$\beta$ $=0.249^{+0.008}_{-0.008},$
$\gamma$ $=-0.555^{+0.018}_{-0.018}.$
The corner plot in Figure B9 demonstrates that all parameters are well-constrained, with physically plausible correlations and no significant degeneracies. This calibration provides a statistically robust foundation for the application of the SHMR-based halo mass estimator to both simulated and observational group catalogues. For further details on the calibration methodology and performance, see Sections 3.2 and 3.2.3.

### v3

@@ Appendix B SHMR Fitting Parameters
The calibration of the sSHMR was performed using a Bayesian MCMC approach, enabling a rigorous exploration of the posterior probability distributions for all model parameters. The adopted functional form for the SHMR is a double power-law,
@@
The resulting posterior distributions and covariances for the four free parameters are shown in Figure B10. The median and the $16^{th}$ and $84^{th}$ percentile for each parameter given in Table B4.
Table B4: Best-fitting of the coefficients for the sSHMR, including their 16th and 84th percentile.
Coefficient $16^{th}$ Percentile Median $84^{th}$ Percentile
AA 46.43 46.944 47.47
$\log M_{A}[M_{\odot}]$ 10.453 10.483 10.515
$\beta$ 0.241 0.249 0.257
$\gamma$ -0.573 -0.555 -0.537
The corner plot in Figure B10 demonstrates that all parameters are well-constrained, with physically plausible correlations and no significant degeneracies. This calibration provides a statistically robust foundation for the application of the SHMR-based halo mass estimator to both simulated and observational group catalogues. For further details on the calibration methodology and performance, see Sections 4.2 and 4.2.3.